# Multimodal Medical Image Binding via Shared Text Embeddings

### v1

@@ 4.5. Evaluations in Few-Shot Manner
Figure 3. Comparison of Few-Shot Performance Across Various Datasets (ODIR200x3, CheXpert, SICAPv2, RAD-ChestCT) With Increasing Shot Numbers.
@@
4.6. Evaluations in Cross-Modal Manner
Table 3. Cross-modal retrieval performance (Recall Top@1, 5, 10) comparison of M3Bind against two different variants of MedBind. Red and blue highlighting indicate the best and second-best performance, respectively.
@@ 4.6. Evaluations in Cross-Modal Manner
4.7. Ablation Study
Ablation Analysis of X-Ray, ECG, Retina, Pathology, and CT Modalities. We conducted an ablation study to evaluate the contributions of X-ray, ECG, retina, pathology and CT images on the ODIR200x31 and MIMIC-PAIR datasets, as summarized in Table 4. As the number of modalities involved in training increases, the model’s performance gradually improves, especially when the CT modality, which is highly relevant to chest X-Ray and ECG, is added. The experimental results demonstrate that each additional modality incrementally improves the quality of the shared space representation.
@@ 4.7. Ablation Study
Table 4. Ablation Study on Individual Contributions of X-Ray, ECG, Retina, Pathology, and CT Modalities: Performance is evaluated on the MIMIC-PAIR datasets.
@@
Table 5. Ablation Study of AMB and SESKD Modules: Performance Evaluation on the ODIR200x3 and MIMIC-PAIR Datasets Using the M3Bind Framework.
M3Bind AMB SESKD ODIR200x31 MIMIC-PAIR
@@
5. Conclusion
In this paper, we introduce M3Bind, a novel multimodal pre-training method for aligning medical imaging representation. Using an anchored text modality, M3Bind harmonizes image representations across five CLIP-like medical models, allowing cross-modal alignment without relying on arbitrary paired data between two medical image modalities. Specifically, our framework employs contrastive learning with MSE loss to align text embedding spaces while preserving the original image-text relationships. M3Bind addresses data imbalance challenges through adaptive modality balancing, ensuring learning across all modalities, and introduces the distillation of knowledge of shared embedding space to consolidate modality-specific text encoders into a unified encoder. The future direction can explore the applications of M3Bind in building a multimodal large langue model and scaling to more available modalities. Our work focuses on cross-modal alignment in medical imaging but only investigates the cross-modal retrieval task between X-Ray and ECG, so another direction is the evaluations on other cross-modal settings.

### v2

@@ 4.4 Evaluations in Few-Shot Manner
Figure 3: Comparison of Few-Shot Performance Across Various Datasets (ODIR200x3, CheXpert, SICAPv2, RAD-ChestCT) With Increasing Shot Numbers.
@@
4.5 Evaluations in Cross-Modal Manner
Table 3: Cross-modal retrieval performance (Recall Top@1, 5, 10) comparison of M3Bind against two different variants of MedBind. Red and blue highlighting indicate the best and second-best performance, respectively.
@@ 4.5 Evaluations in Cross-Modal Manner
4.6 Ablation Study
Ablation Analysis of X-Ray, ECG, Retina, Pathology, and CT Modalities. We conducted an ablation study to evaluate the contributions of X-ray, ECG, retina, pathology and CT images on the ODIR200x3LABEL:1 and MIMIC-PAIR datasets, as summarized in Table 4. As the number of modalities involved in training increases, the model’s performance gradually improves, especially when the CT modality, which is highly relevant to chest X-Ray and ECG, is added. The experimental results demonstrate that each additional modality incrementally improves the quality of the shared space representation.
@@ 4.6 Ablation Study
Table 4: Ablation Study on Individual Contributions of X-Ray, ECG, Retina, Pathology, and CT Modalities: Performance is evaluated on the MIMIC-PAIR datasets.
@@
Table 5: Ablation Study of AMB and SESKD Modules: Performance Evaluation on the ODIR200x3 and MIMIC-PAIR Datasets Using the M3Bind Framework.
M3Bind AMB SESKD ODIR200x3LABEL:1 MIMIC-PAIR
@@
5 Conclusion
In this paper, we introduce M3Bind, a novel multimodal pre-training method for aligning medical imaging representation. Using an anchored text modality, M3Bind harmonizes image representations across five CLIP-like medical models, allowing cross-modal alignment without relying on arbitrary paired data between two medical image modalities. Experimental results show that our framework not only achieves implicit alignment across different image modalities, but also brings performance improvements to the original models on image-text tasks. The flexibility of our framework allows it to benefit from stronger CLIP-like models—if we select more powerful base models, our framework can accordingly achieve even better performance, potentially reaching state-of-the-art (SOTA) results.